OpenStudy (anonymous):

A 130.0−mL sample of a solution that is 3.0×10−3M in AgNO3 is mixed with a 225.0−mL sample of a solution that is 0.14M in NaCN. After the solution reaches equilibrium, what concentration of Ag+(aq) remains?

OpenStudy (abmon98):

Number of Moles=Concentration*Volume 130/1000*3.0*10^-3=0.00039 moles of AgNO3 Number of Moles=Concentration*Volume 225/1000*0.14=0.0315 moles of NaCN Sodium is reactive enough to displace Ag+ to produce NaNO3 when 0.0003375 moles of Ag+ is mixed with 0.0322 moles of CN- we say that essentially all of the 0.00039 moles of Ag+ is converted to AgCN the 0.00039 moles of Ag+ has been diluted to a total volume of 225+130=355 ml: Number of Moles=Concentration*Volume 0.00039=355/1000*Concentration C=0.00109859154 mol/dm^3 We are Dividing by 1000 to change the units from ml to dm^3